OpenStudy (anonymous):

find the value of sqrt 8/9

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large \sqrt{8 \over 9} = {\sqrt8 \over \sqrt9} }\)

Parth (parthkohli):

Simplify \(\sqrt8\)

OpenStudy (unklerhaukus):

\[\sqrt8=\sqrt{2\times2\times 2}=\sqrt{2^2\times2}=\cdots\] \[\sqrt9=\sqrt{3\times3}=\sqrt{3^2}=\cdots\]

OpenStudy (anonymous):

1 1/3 ?

OpenStudy (anonymous):